Nicoles Getaway Spa (NGS) has been so successful that Nicole has decided to expand her spa by selling merchandise. She sellsRequired: 1. Prepare journal entries for each transaction. (If no entry is required for a transaction/event, select No JournView transaction list 1 NGS sold 10 items of merchandise to Salon World on account at a selling price of $1,000 (total); term6 Record the entry for cost of item returned. Collected payment from Salon World, fully paying off the account balance. 8 Sol2. Calculate the amount of Net Sales and Cost of Goods Sold for the transactions listed above. What is Nicoles Getaway Spas

Please note that there are ten transactions in part one. Please label all the ten transactions in the journal entry worksheets respectively. Thank you.
Answer #1

1. Journal Entries

$ $
1. 2/8 Salon Word a/c Dr 1000
To Sales a/c 1000
(sales made)
2. 2/8 Cost of Goods Sold Dr 580
To Inventory a/c 580
(Goods sold)
3. 3/8 Cosmetics 4 You a/c Dr 790
To Sales a/c 790
(sales made)
4. 3/8 Cost of Goods Sold Dr 365
To Inventory a/c 365
(Goods sold)
5. 6/8 Sales (Returns) a/c Dr (790/5) 158
To Cosmetics 4 You 158
(Goods returned)
6. 6/8 Inventory a/c Dr(365/5) 73
To Cost of goods sold a/c 73
(Goods returned to Inventory)
7. 10/8 Cash a/c Dr 990
Sales Discount a/c Dr (1000*1%) 10
To Salon World a/c 1000
(Cash received within the discount period)
8. 20/8 Meghan Witzel a/c Dr 265
To Sales a/c 265
(sales made)
9. 20/8 Cost of Goods Sold a/c Dr 79
To Inventory a/c 79
(goods sold)
10. 22/8 Cash a/c Dr (790*4/5) 632
To Cosmetics 4 You 632
(Cash received after discount period)

2.based on above jounral entries we get,

  • Net Sales: (1000+790-158 (returns) +265) =1897
  • Cost of Goods sold:(580+365-73(returns)+79) = 951
  • Gross Profit (%): (Sales-COGS)/Sales*100 ie., (1897-951)/1897*100 = 49.86%